# Forgetting Order of Continual Learning: Examples That are Learned First are Forgotten Last

### v1

@@ Appendix A Class incremental learning
Figure 9: Goldilocks vs. uniform sampling with various continual learning methods, in class incremental settings. Each bar shows the mean final test accuracy across all tasks for each method, with a uniform sampled buffer (traditionally done) in blue, and Goldilocks in orange. Error bars denote the standard error.
@@
Figure 10: Extendeing Fig.2 to CIFAR-100-2, Cifar-10-2 and a subset of TinyImagenet-2. (top) The first task’s binary epoch-wise classification matrices $M$ for the test data of each dataset. Rows denote examples, and columns denote epochs, indicating correct classifications by a majority of the networks. Rows are sorted by learning speed. Faster-learned examples from the first task are less likely to be forgotten at the end of the second task. (bottom) The mean learning speed of examples in the first task vs. the percentage of networks that remember them at the end of the second task. Networks forget more examples learned slowly across all 3 datasets.
@@
Figure 12: Comparing the extended epoch-wise classification matrix for different architectures, rows represent examples, and columns represent epochs, indicating correct classifications by a majority of 10 networks. In (a), we train ResNet-18. In (b), we train ResNet-18 with both width and depth reduced by half. In (c), we train ResNet-18 with both width and depth reduced by a factor of four. In all cases, rows are sorted by learning speed. Examples learned quickly in the first task are less likely to be forgotten after the second task.
@@
Figure 13: Comparing different replay buffer compositions for various continual learning methods, with a buffer of $1\%$. The experimental setup replicates the one done in Fig. 3.2. Like the experience replay case, focusing on examples learned midway through the learning process is most beneficial. Due to the smaller buffer, removing more easier examples is better.
Appendix B Correlations between learning speed and catastrophic forgetting in different datasets
@@
Other Datasets. We first explore other datasets using the experimental setup detailed in Fig. 2. We extend our analysis to CIFAR-10-2, CIFAR-100-2, and a subset of TinyImagenet-2 (comprising the initial 40 classes). The extended epoch-wise classification matrices for each task and the correlation between learning speed and the percentage of networks remembering each example are plotted in Fig. 10. Consistently across all three datasets, we observe a robust correlation between learning speed and the likelihood of example retention during continual training. Notably, faster-learned examples exhibit lower rates of catastrophic forgetting, as evidenced by both quantitative correlations and visual inspection of the extended epoch-wise classification matrices.
@@
Other Architectures. We extend the results from Fig. 2 to architectures beyond ResNet-18 by creating two new, smaller variants: Small-ResNet and Tiny-ResNet. These are formed by reducing both the width and depth of ResNet-18 by a factor of 2 and 4 respectively. We replicate the experimental setup from Fig.2, training 10 models of each architecture on CIFAR-100-2 without a replay buffer, and storing the extended epoch-wise classification matrix. Each matrix row is sorted by the learning speed of the examples and plotted in Fig. 12. All matrices depict the classification of the test dataset. In all architectures, learning speed is highly correlated with catastrophic forgetting: networks forget more examples learned later in training while retaining an almost perfect recollection of those learned early on. Additionally, stronger architectures, similar to larger buffers, can remember slower-to-learn examples. This reinforces the motivation for Goldilocks, as it shows that the examples selected by Goldilocks are those the model could learn independently if it had a stronger architecture.
@@
Other continual learning algorithms. In §3.4, we demonstrated the benefits of non-uniform buffer compositions for various continual learning methods. We extend these findings, initially shown for GEM in Fig. 6, to additional algorithms including DER, DER++, GEM, A-GEM, RPC, and X-DER. For each algorithm, we trained 10 networks with different buffer compositions by varying the $quick$ and $slow$ hyperparameters of Goldilocks. For each algorithm, we trained 10 networks on different buffer compositions, achieved by varying the $quick$ and $slow$ hyperparameters of Goldilocks. In all cases, we used the hyperparameters and architectures suggested in the original works. We used a buffer of size $500$ ($1\%$ of the data), due to its popularity in previous works. Consistent with the GEM results, we found that replay buffers focused on examples learned mid-way through the training process were most beneficial. Additionally, in all cases, removing slower-to-learn examples proved advantageous, given the small buffer size. These results can be found in Fig. 13.
Different number of tasks.
In most experiments presented in the paper, we focused on continual learning with 2 tasks. When considering a larger number of tasks, additional hyperparameters emerge, as the $quick$ and $slow$ parameters apply to each intermediate task. These can be tuned similarly to the two-task case, either through grid-search with a rotation task or using simple heuristics. To avoid excessive hyperparameter tuning when training on more than two tasks, we tuned the $quick$ and $slow$ values for the first task and kept them constant for all subsequent tasks. Results from training on CIFAR-10-5, CIFAR-100-5, and TinyImageNet-10 are shown in Fig. 17. We observe that Goldilocks improves over uniform sampling in these cases as well. Notably, further tuning of all Goldilocks hyperparameters could yield even better results.
Appendix C standard errors

### v2

@@ 3.2 SBS Improves Various CL Algorithms
Figure 6: SBS vs. uniform sampling with various continual learning methods. Bars show the mean final test accuracy across all tasks, with random sampling in blue, and SBS in orange. Error bars denote the standard error. (a) CIFAR-100-5 (b) TinyImageNet-10, with a buffer of $500$ examples. SBS improves performance across all methods and datasets.
Many replay-based CL algorithms rely on uniform sampling for their replay buffers. Here, we evaluate the impact of replacing uniform sampling with SBS across such algorithms, including DER, DER++ (Buzzega et al., 2020), X-DER (Boschini et al., 2022), AGEM (Chaudhry et al., 2019), ER (Rolnick et al., 2019), GEM (Lopez-Paz & Ranzato, 2017), RPC (Pernici et al., 2021), and FDR (Titsias et al., 2020). Our results show that SBS consistently improves the performance of all tested continual algorithms, advancing the state of the art in multiple image classification tasks.
We compared the mean final test accuracy across tasks for each method trained with either uniform sampling (as in the original works) or SBS. While not all methods perform equally well in every scenario, SBS significantly improves accuracy compared to random sampling in all cases. We focused on small buffer sizes, as these are the primary use cases addressed in the original works.
Fig. 6 shows results on CIFAR-10-5, CIFAR-100-5, and TinyImageNet-10 with a fixed buffer size of $500$. Each task was trained for 50 epochs. As the buffer size is fixed, examples were removed starting from the second task to accommodate new data. For SBS, examples were removed at random, while uniform sampling employed either random or reservoir sampling, following the strategies outlined in each method’s original implementation. The $q$ and $s$ hyperparameters for SBS were determined using the RotNet auxiliary task, as detailed in §3.1.
Table 1: Comparison of replay buffer sampling methods. Each entry represents the mean test accuracy across all tasks for 10 networks trained with experience replay using different sampling methods. The highest accuracy in each scenario is in bold. For visualization, standard errors are reported separately in Table E. While certain methods, such as Herding (small buffers) and GSS (large buffers), perform well in specific cases, they fail to consistently outperform random sampling across all scenarios. In contrast, SBS consistently achieves the highest accuracy across all scenarios.
CIFAR-100-20 CIFAR-10-5 TinyImageN-2
Buffer $1k$ $10k$ $1k$ $10k$ $1k$ $10k$
Random 51.75 71.25 79.48 82.4 49.81 61.48
Max Ent 48.3 69.83 77.28 81.72 44.49 57.88
IPM 49.91 71.7 79.3 80.57 48.58 61.97
GSS 48.13 72.9 76.71 84.13 49.36 62.89
Herding 54.0 73.51 81.8 81.03 51.17 62.44
LARS 51.82 71.41 79.17 84.93 50.22 62.78
SBS-fix 54.65 73.72 82.24 86.15 52.15 63.12
SBS 55.43 74.80 83.59 89.17 52.15 63.16
When training with multiple tasks, different $q$ and $s$ hyperparameters could be used for each task. However, to reduce computational cost and for simplicity, we apply the same $q$ and $s$ values across all tasks in multi-task scenarios. While choosing separate $q$ and $s$ for each task would likely improve performance further, we observe that SBS’s performance boost is especially pronounced in multi-task settings (see App. C.5), indicating its particular effectiveness in such scenarios.
3.3 SBS Vs. Other Replay Buffer Sampling Methods
While numerous replay buffer sampling methods have been proposed over the years, uniform sampling remains prevalent in practical implementation within CL algorithms. This preference may stem from the specific settings or algorithmic modifications required for other sampling methods to work well. These methods do not consistently improve different CL algorithms across datasets and buffer sizes. In contrast, SBS improves a wide range of continual learning algorithms across various buffer sizes and datasets.
We explored various replay buffer sampling strategies, including uniform sampling, max entropy, IPM (Zaeemzadeh et al., 2019), GSS (Aljundi et al., 2019), Herding (Rebuffi et al., 2017), and LARS (Buzzega et al., 2021). We trained $10$ networks on TinyImageNet-2, CIFAR-100-20, and CIFAR-10-5 datasets with buffer sizes of $1k$ and $10k$ using experience replay. Table 3.2 presents the mean test accuracy across all tasks for each case. Additional results on CIFAR-10-2 and CIFAR-100-2 can be found in App. C.5.
SBS consistently outperforms all other sampling strategies across all evaluated scenarios. In contrast, alternative methods show effectiveness only in specific settings. For example, Herding – favoring examples whose features are close to the class mean – surpasses uniform sampling at smaller buffer sizes but underperforms with larger ones. This likely reflects its bias toward typical examples, which are learned early and thus more useful when the replay buffer is small. Conversely, GSS, which prioritizes examples with large gradient distances, excels with larger buffers but lags with smaller ones, as its preference for slowly learned examples becomes more beneficial when a larger replay buffer is available. Table 3.2 also includes SBS-fix, a variant of SBS that uses default hyperparameters ($q=s=20\%$, as recommended in §3.1) without fine-tuning. While SBS-fix underperforms SBS, it still consistently outperforms other sampling methods, indicating that extensive hyperparameter tuning is often unnecessary.
3.4 Limitations
One limitation of SBS is that the learning speed score is based on the number of training epochs. With a small number of epochs, the score becomes overly discrete, which reduces the precision of the evaluation. However, as shown in App. D and Fig. 20, SBS remains effective even with a small number of epochs, as long as the network converges, although with increased noise. Nonetheless, many continual learning scenarios emphasize stream-like settings, which typically involve only a single epoch. These scenarios are not well-supported by SBS, as the learning speed fails to adequately capture how different examples are learned over time. Alternative methods for approximating example complexity are needed, which we leave as future work.
3.5 Discussion
This paper takes an observational approach to catastrophic forgetting, identifying a link between the speed at which examples are learned and their likelihood of being forgotten – examples learned more quickly are less prone to forgetting when new data is introduced. SBS was proposed as a simple sampling function that demonstrates the practical utility of this observation. By modifying the contents of the replay buffer, SBS encourages algorithms to focus on examples learned at varying speeds. Indeed, integrating SBS into a wide range of continual learning algorithms across diverse settings yields significant improvements, suggesting that this link is an important factor in mitigating catastrophic forgetting. This insight opens avenues for future work, such as designing more sophisticated sampling strategies or incorporating learning speed more directly into continual learning methods.
The simplicity of SBS also brings practical advantages: it is computationally efficient, easy to implement, and compatible with any buffer-based continual learning approach. As continual learning methods evolve, SBS remains relevant due to its ease of integration and broad applicability.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none of which we feel must be specifically highlighted here.
Acknowledgment
This work was supported by ERC-2020-AdG grant 101021347, KeepOnLearning.
Appendix
Appendix A Implementation Details
Architectures and hyper-parameters. In our experiments, unless stated otherwise, we trained ResNet-18 for $E=100$ epochs per task. We employed a base learning rate of $0.1$ with a cosine scheduler, SGD optimizer, momentum of $0.9$, and weight decay of $0.0005$. All networks were trained on NVIDIA TITAN X. These hyperparameters were chosen arbitrarily based on their performance during joint dataset training and were consistent across all experiments. We anticipate consistent qualitative results despite potential variations in these hyperparameters. When introducing a replay buffer, we employed an experience replay strategy (Rolnick et al., 2019), alternating batches of data from the new task and the replay buffer. When integrating with other continual learning algorithms, we evaluated all methods within the framework of (Boschini et al., 2022; Buzzega et al., 2020), using the hyperparameters suggested in the original papers, changing only the replay buffer sampling to SBS, keeping the rest of the method intact.
Appendix B Class Incremental Learning
There are two primary paradigms in continual learning: class incremental learning (CIL) and task incremental learning (TIL). The distinction lies in how tasks are handled. In TIL, task identities are known during inference, enabling the use of specialized classifiers within each task, whereas in CIL, the model must infer both the task and class identity without prior knowledge, often leading to lower performance. These paradigms address different real-world scenarios, and methods optimized for one sometimes underperform on the other, necessitating separate evaluations in continual learning research.
Our work examines the relationship between the speed at which examples are learned and their susceptibility to forgetting in continual learning. Despite the differences between CIL and TIL, our study reveals that the same qualitative results and the same conclusions can be made in both cases. Therefore, to avoid repetition in the main text, we focus there on TIL only, as it isolates forgetting at the model level rather than conflating it with final-layer interference. To complement this, we reproduce all key experiments under CIL settings in this appendix, highlighting similar conclusions despite the inherent challenges of CIL.
Figs. 7, 8, 9, 10, 17 present CIL counterparts to the main text’s results, Figs. 2, 3, 4, 6, 16 respectively. While CIL results exhibit lower overall performance, the trends remain consistent: examples learned faster are less likely to be forgotten, and the correlations between the buffer size and forgetting and the correlation between the learning speed and forgetting are strong. Notably, optimal buffer composition in CIL scenarios (Figs. 9, 17) slightly favors keeping more quickly learned examples compared to TIL, reflecting the added complexity of task inference in CIL.
Fig. 10 reproduces results from Fig. 6, demonstrating the impact SBS sampling strategy across CIL datasets and algorithms. SBS consistently yields significant performance gains, reaffirming its effectiveness across diverse continual learning challenges.
Finally, Table B shows the results of the experiments in Table 3.2 in the CIL settings, showing similar qualitative behavior.
Table 2: The same table as Table 3.2, when doing class-incremental learning instead of task-incremental. Each entry represents the mean test accuracy across all tasks for 10 networks trained with experience replay using different sampling methods. The highest accuracy in each scenario is in bold.
CIFAR-100-20 CIFAR-10-5 TinyImageN-2
Buffer $1k$ $10k$ $1k$ $10k$ $1k$ $10k$
Random 42.9 48.5 57.74 79.47 9.99 27.4
Max Ent 41.76 47.82 57.11 69.13 9.47 27.2
IPM 42.94 49.51 57.6 80.87 9.96 27.48
GSS 41.63 49.12 57.73 71.27 9.19 27.73
Herding 44.65 48.47 58.02 79.02 9.38 28.43
LARS 43.55 49.73 57.32 79.96 10.21 28.1
SBS-fix 44.87 50.39 58.89 81.16 12.78 29.37
SBS 45.64 53.71 61.13 82.22 12.78 29.41
(a) Test accuracy
(b) $M$ train
(c) $M$ test
(d) Speed vs. forgetting
[…]
(a) Small buffer
(b) Medium buffer
(c) Large buffer
(d) Buffer size vs. speed
Figure 8: Repeating Fig. 3 for class incremental learning. Impact of the buffer size on forgetting dynamics. (a-c) The first task’s binary matrices $M$, when adding replay buffers of varying buffer sizes. (d) The mean learning speed of remembered examples of models with different buffer sizes. Models with bigger replay buffers remember slower-to-learn examples.
(a) CIFAR-100-2, small buffer
(b) CIFAR-100-2, big buffer
Figure 9: Repeating Fig. 4 for class incremental learning. Comparison of buffer compositions for various buffer sizes. (a) CIFAR-100-2 with a small $1k$ buffer. (b) CIFAR-100-2 with a large $10k$ buffer. Note that since class incremental learning is harder than task incremental learning, it is beneficial to remove fewer "quick" examples in this case.
(a) CIFAR-100-5
(b) CIFAR-10-5
(c) TinyImageNet-10
Figure 10: SBS vs. uniform sampling with various continual learning methods, in class-incremental settings. Each bar shows the mean final test accuracy across all tasks for each method, with a uniform sampled buffer (traditionally done) in blue, and SBS in orange. Error bars denote the standard error.
@@ Appendix B Class Incremental Learning
Figure 11: Extendeing Fig. 2 to CIFAR-100-2, Cifar-10-2 and a subset of TinyImageNet-2. (top) The first task’s binary epoch-wise classification matrices $M$ for the test data of each dataset. The y-axis denotes examples, and the x-axis denotes epochs, indicating if an example is correctly classified by the model at the given epoch. The order in which examples appear is sorted by learning speed. Faster-learned examples from the first task are less likely to be forgotten at the end of the second task. (bottom) The mean learning speed of examples in the first task vs. the percentage of networks that remember them at the end of the second task. Networks forget more examples learned slowly across all 3 datasets.
@@
Figure 13: Comparing the extended epoch-wise classification matrix for different architectures, the y-axis represents examples, and the x-axis represents epochs, indicating if an example is correctly classified by the model in a given epoch. In (a), we train ResNet-18. In (b), we train ResNet-18 with both width and depth reduced by half. In (c), we train ResNet-18 with both width and depth reduced by a factor of four. In all cases, the order examples appear is sorted by learning speed. Examples learned quickly in the first task are less likely to be forgotten after the second task.
@@
Figure 14: Comparing different replay buffer compositions for various continual learning methods, with a buffer of size $500$. The experimental setup replicates the one done in Fig. 4. Like the experience replay case, focusing on examples learned midway through the learning process is most beneficial. Due to the smaller buffer, removing slower examples is better.
Appendix C Correlations Between Learning Speed and Catastrophic Forgetting in Different Datasets
@@
C.1 Other Datasets
We first explore other datasets using the experimental setup detailed in Fig. 2. We extend our analysis to CIFAR-10-2, CIFAR-100-2, and a subset of TinyImageNet-2 (comprising the initial 40 classes). The extended epoch-wise classification matrices for each task and the correlation between learning speed and the percentage of networks remembering each example are plotted in Fig. 11. Consistently across all three datasets, we observe a robust correlation between learning speed and the likelihood of example retention during continual training. Notably, faster-learned examples exhibit lower rates of catastrophic forgetting, as evidenced by both quantitative correlations and visual inspection of the extended epoch-wise classification matrices.
@@ C.1 Other Datasets
C.2 Classification Matrices of Different Datasets
In Section 2, we analyze epoch-wise classification matrices, which indicate whether each example was classified correctly or incorrectly at each epoch during training. Figs. 2 and 3 plot these matrices for CIFAR-100-2 and CIFAR-10-2, revealing a clear correlation between learning speed and catastrophic forgetting: examples learned quickly are less likely to be forgotten during task switches.
In Fig. 15, we extend this analysis to the classification matrices of the first task of TinyImageNet-2, CIFAR-100-20, and CIFAR-10-5. For these datasets, which involve more than two tasks, the x-axis includes multiple task switches, marked by vertical dashed black lines. The same correlation is observed and becomes even more pronounced: examples from the first task are increasingly forgotten as additional tasks are introduced. With each new task, the newly forgotten examples are increasingly those learned earlier in training. Notably, examples learned fastest remain robust to forgetting, even after a long sequence of tasks, while those learned more slowly are more vulnerable.
(a) TinyImageNet-2
(b) CIFAR-10-5
(c) CIFAR-100-20
Figure 15: Repeating Fig. 2(c) with different datasets and task counts: (a) TinyImageNet with 2 tasks, (b) CIFAR-10 with 5 tasks, and (c) CIFAR-100 with 20 tasks. Dashed black lines indicate the introduction of a new task. Consistent with Fig. 2(c), examples learned more slowly are more prone to forgetting after task transitions. As the number of tasks increases, even examples learned relatively quickly become susceptible to forgetting.
C.3 Other Architectures
@@
We extend the results from Fig. 2 to architectures beyond ResNet-18 by creating two new, smaller variants: Small-ResNet and Tiny-ResNet, which can be found in Fig. 13. These are formed by reducing both the width and depth of ResNet-18 by a factor of 2 and 4 respectively. We replicate the experimental setup from Fig. 2, training 10 models of each architecture on CIFAR-100-2 without a replay buffer, and storing the extended epoch-wise classification matrix. The order examples appear in the matrix is sorted by the learning speed of the examples. All matrices depict the classification of the test dataset. In all architectures, learning speed is highly correlated with catastrophic forgetting: networks forget more examples learned later in training while retaining an almost perfect recollection of those learned early on. Additionally, stronger architectures, similar to larger buffers, can remember slower-to-learn examples. This reinforces the motivation for SBS, as it shows that the examples selected by SBS are those the model could learn independently if it had a stronger architecture.
C.4 Other Continual Learning Algorithms
@@
In §3.2, we demonstrated the benefits of non-uniform buffer compositions for various continual learning methods. We extend these findings to additional algorithms including DER, DER++, GEM, A-GEM, RPC, and X-DER. For each algorithm, we trained 10 networks with different buffer compositions by varying the $quick$ and $slow$ hyperparameters of SBS. For each algorithm, we trained 10 networks on different buffer compositions, achieved by varying the $quick$ and $slow$ hyperparameters of SBS. In all cases, we used the hyperparameters and architectures suggested in the original works. We used a buffer of size $500$, due to its popularity in previous works. Consistent with the GEM results, we found that replay buffers focused on examples learned mid-way through the training process were most beneficial. Additionally, in all cases, removing slower-to-learn examples proved advantageous, given the small buffer size. These results can be found in Fig. 14.
C.5 Two Tasks Vs. Multi-Task
In the behavioral analysis, we focused on a two-task setting as it provides a simple and controlled environment. However, SBS was introduced in multi-task scenarios, which are standard in the CL literature and allow for more meaningful comparisons. For completeness, we now extend our SBS results to the two-task setting and the behavioral results to multi-task scenarios, aligning them with each other.
Fig.4 expands on Fig.6 by showcasing a broader range of SBS hyperparameters, hence including the results of SBS on these datasets. Using the RotNet auxiliary task to pick the hyperparameter in those cases obtains the hyperparameters that maximize performance across all three cases.
Table C.5 extends Table 3.2 with results on CIFAR-10-2 and CIFAR-100-2, showing trends consistent with the multi-task setting. Additionally, Figs.16 and 17 extend Fig.4 to CIFAR-100-20 and CIFAR-10-5 (buffer size 5k) in both class- and task-incremental settings, showing similar qualitative trends.
Notably, while SBS improves performance in both two- and multi-task scenarios, its impact is even more pronounced in multi-task settings, further highlighting its practical value.
Table 3: Extended results from Table 3.2 for CIFAR-10-2 and CIFAR-100-2. For convenience, the original results from Table 3.2 are included. Standard errors are reported separately in Table E.
CIFAR-10-2 CIFAR-100-2 CIFAR-100-20 CIFAR-10-5 TinyImageNet-2
Buffer size $1k$ $10k$ $1k$ $10k$ $1k$ $10k$ $1k$ $10k$ $1k$ $10k$
Random 87.67 95.03 69.03 79.24 51.75 71.25 79.48 82.4 49.81 61.48
Max entropy 84.44 94.63 64.27 77.91 48.3 69.83 77.28 81.72 44.49 57.88
IPM 85.9 94.76 64.16 75.73 49.91 71.7 79.3 80.57 48.58 61.97
GSS 85.28 95.55 64.15 80.11 48.13 72.9 76.71 84.13 49.36 62.89
Herding 88.41 94.53 71.07 78.97 54.0 73.51 81.8 81.03 51.17 62.44
LARS 87.43 95.43 69.53 79.92 51.82 71.41 79.17 84.93 50.22 62.78
SBS-fix 88.11 95.56 70.29 80.2 54.65 73.72 82.24 86.15 52.15 63.12
SBS 88.99 96.03 71.71 80.62 55.43 74.80 83.59 89.17 52.15 63.16
(a) CIFAR-100-20
(b) CIFAR-10-5
Figure 16: Comparison of buffer compositions in multi-task training scenarios: (a) CIFAR-100-20, (b) CIFAR-10-5, both trained with a 5k buffer. Similarly to the 2-task case, a wide range of buffer compositions significantly improves performance, helping to mitigate catastrophic forgetting. The performance gains of Goldilocks tend to get bigger in the multi-task case.
(a) CIFAR-100-20
(b) CIFAR-10-5
Figure 17: Repeating Fig. 16 for class incremental learning. (a) CIFAR-100-20, (b) CIFAR-10-5, both with a 5k buffer. Similar to the 2-task case, a wide range of buffer compositions significantly improves performance, helping to mitigate catastrophic forgetting.
Appendix D The Effects of Different Learning Hyper-Parameters
In this section, we investigate how varying learning hyperparameters influences the optimal composition of examples in the replay buffer. We replicate the experiment shown in Fig. 4(a), training networks on CIFAR-100-2 with a buffer size of 1000, while modifying factors such as network architecture, optimizer choice, learning rate, regularization strength, and the number of training epochs.
Across all configurations, we observe consistent qualitative trends similar to those in Fig. 4(a): a wide range of buffer compositions significantly improves performance and alleviates catastrophic forgetting. These findings suggest that the conclusions drawn in the main paper are robust and generalizable, extending to continual learning scenarios under different hyperparameter settings.
Optimizers. In Fig. 18, we compare the performance of different buffer compositions when training ResNet-18 on CIFAR-100-2 using three optimizers: SGD, Adam, and Adagrad. For SGD, we used a momentum of 0.9, a weight decay of $0.0005$, and a learning rate of $0.1$, as is common for CIFAR-100 training. For Adagrad, we employed a learning rate of $0.01$ and a weight decay of $0.0001$. For Adam, we used a learning rate of $0.001$, a weight decay of $0.0001$, and $(0.9,0.999)$ betas. These hyperparameters were tuned to optimize ResNet-18’s performance on CIFAR-100 without considering continual learning, as the optimal settings vary by optimizer due to their differing update mechanisms.
Our results indicate that similar buffer compositions consistently enhance performance across all three optimizers. This consistency suggests that the benefits of our analysis are robust to the choice of optimizer, further supporting the generalizability of our findings.
(a) SGD
(b) Adam
(c) AdaGrad
Figure 18: Repeating Fig. 4(a) using different optimizers: (a) SGD, (b) Adam, and (c) AdaGrad. Across all optimizers, a wide range of buffer compositions consistently enhances performance, yielding similar qualitative results to the experiment in the main text.
Architectures. In Fig. 19(a), we evaluate the impact of buffer composition on the performance of three different architectures: ResNet-18, VGG-16, and a smaller version of ResNet, tiny-ResNet. All models were trained on CIFAR-100-2 under identical conditions. Our results show that similar buffer compositions consistently improve learning performance across these diverse architectures. This finding suggests that our analysis is not specific to any single architecture.
(a) VGG-16
(b) No regularization
(c) Different CIFAR Split
Figure 19: Repeating Fig. 4(a) under different learning settings: (a) replacing the ResNet-18 architecture with VGG-16, (b) removing weight decay from ResNet-18, and (c) using a random class-to-task split of CIFAR-100-2. In all cases, the results remain qualitatively similar, highlighting the robustness of the analysis.
Training epochs. In Fig. 20, we compare the performance of models trained for different numbers of epochs per task, all using a cosine learning rate scheduler to ensure exposure to both large and small learning rates throughout training. While reducing the number of epochs leads to a general drop in performance, we observe that similar buffer compositions consistently yield strong results, even under shorter training schedules. It is worth noting that the accuracy of determining the learning speed of each example may decrease with fewer epochs, as there is less time for the model to adapt. Nonetheless, the qualitative trends remain robust, with specific buffer compositions consistently outperforming or underperforming the random baseline across all tested epoch counts.
(a) 20 epochs
(b) 50 epochs
(c) 100 epochs
Figure 20: Repeating Fig. 4(a) with varying training epochs: (a) 20 epochs, (b) 50 epochs, and (c) 100 epochs per task. Despite shorter training durations, the qualitative results remain consistent, with a wide range of buffer compositions significantly improving performance.
Learning Rates. We examine the impact of learning rates on model performance using SGD with rates of $0.05$ and $0.2$ (Figs. 21(a), 21(b)). While the final accuracies differ across these configurations, the qualitative behavior remains consistent: the same buffer composition consistently achieves strong results.
Fine-Tuning. In multi-task training, it is common to reduce the learning rate for subsequent tasks to fine-tune the network on the next task. To investigate this, we replicate the experiment in Fig. 4(a), lowering the learning rate for the second task by a factor of $10$. The results align closely with those in Fig. 4(a), indicating that our findings are robust and not solely attributable to the specific learning rates used in the original experiments. These results can be found in Fig. 21(c).
(a) Smaller learning rate
(b) Larger learning rate
(c) Finetune
Figure 21: Repeating Fig. 4(a) with varying learning rates: (a) doubling the learning rate, (b) halving the learning rate, and (c) keeping the first task’s learning rate unchanged while reducing the second task’s by a factor of 10. These scenarios, common in continual learning research, yield consistent qualitative results, indicating that our analysis is not dependent on a specific learning rate.
Regularization. All ResNet-18 experiments in the main paper were conducted with a small weight decay of $0.0005$. To assess the impact of this regularization, we replicate the experiment from Fig. 4(a) in Fig. 19(b) where we remove the weight decay entirely. The results remain qualitatively consistent, indicating that the conclusions from our analysis are not sensitive to this specific regularization choice.
Different splits of CIFAR. We repeated the experiment from Fig. 4(a) using a different split of CIFAR-100, where classes were randomly assigned to tasks. The results, shown in Fig. 19(c), remain consistent with those in Fig. 4(a). Importantly, all task splits in our experiments were across the paper chosen arbitrarily, ensuring unbiased evaluations.
Appendix E Standard Errors